# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ D.1 Consistency under inner product latent space models
  • $\mathcal{E}_{0}$: Event where $\|x_{i}\|_{2}\leq C_{R}\log n=:R^{\prime}(n)$ for all $i=1..n$. $P(\mathcal{E}_{0})\geq 1-n^{-(k+1)}$.
  • $\mathcal{E}_{1}$: Event where $\|\nabla L(\beta^{*})\|_{2}\leq C_{1}\sqrt{(\log n)/n}$. From Lemma D.1, $P(\mathcal{E}_{1})\geq 1-c_{1}n^{-(k+1)}$.
  • $\mathcal{E}_{2}$: $\nabla^{2}L(\beta)\succeq\mu/4I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. From Lemma D.2, $P(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$.
  • $\mathcal{E}_{3}$: $\nabla^{2}\widetilde{L}(\beta)\succeq\mu/8I=:\tilde{\mu}I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$.
@@
$\bigl{\|}\nabla^{2}\widetilde{L}(\beta)-\nabla^{2}L(\beta)\bigr{\|}_{\rm op}\;\leq\;K\,\delta,$
where $K$ depends only on the $\|x_{i}\|$‐bounds (and thus is $O((\log n)^{2})$ on $\mathcal{E}_{0}$). Hence
$\nabla^{2}\widetilde{L}(\beta)\;\succeq\;\nabla^{2}L(\beta)\;-\;K\delta\,I\;\succeq\;\frac{\mu}{4}\,I\;-\;K\delta\,I.$
@@
$\nabla^{2}\widetilde{L}(\beta)\;\succeq\;\frac{\mu}{8}\,I\quad\mbox{ for }\forall\|\beta-\beta^{*}\|\leq r_{1},$
@@
$\mathcal{E}_{2}\;\subset\;\mathcal{E}_{3}\quad\Longrightarrow\quad P(\mathcal{E}_{3})\;\geq\;P(\mathcal{E}_{2}).$
Let $\mathcal{E}=\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$, and we have $P(\mathcal{E})\geq 1-c_{5}n^{-(k+1)}$.
Next, we work under the conditions from $\mathcal{E}$. We want to use Lemma F.3 to show the unique solution. Set $\phi(\beta)=\nabla\widetilde{L}(\beta)$, $x^{*}=\beta^{*}$, $y^{*}=\nabla\widetilde{L}(\beta^{*})$, $y=0$, $\delta_{1}=r_{1}$.
@@
  • Lower bound: For $\|{\beta-\beta^{*}}\|=r_{1}$. Let $H_{avg}(\beta)=\int_{0}^{1}\nabla^{2}\widetilde{L}(\beta^{*}+t(\beta-\beta^{*}))dt$. The segment $\beta^{*},\beta]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}$, $\nabla^{2}\widetilde{L}(\cdot)\succeq\tilde{\mu}I$ in $B_{r_{1}}(\beta^{*})$, thus $H_{avg}(\beta)\succeq\tilde{\mu}I$. Then, $\|{\phi(\beta)-y^{*}}\|=\|{H_{avg}(\beta)(\beta-\beta^{*})}\|\geq\lambda_{\min}(H_{avg}(\beta))\|{\beta-\beta^{*}}\|\geq\tilde{\mu}r_{1}$. Set $\rho=\tilde{\mu}r_{1}$. The condition $\min_{\|{\beta-\beta^{*}}\|=r_{1}}\|{\phi(\beta)-y^{*}}\|\geq\rho$ holds on $\mathcal{E}_{3}$.
  • Upper bound: We need to show $\|{y-y^{*}}\|=\|{0-\nabla\widetilde{L}(\beta^{*})}\|=\|{\nabla\widetilde{L}(\beta^{*})}\|\leq\rho=\tilde{\mu}r_{1}$. We decompose the gradient using the triangle inequality:
    $\|{\nabla\widetilde{L}(\beta^{*})}\|_{2}$ $=\|{\nabla L(\beta^{*})+\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}$
    $\leq\|{\nabla L(\beta^{*})}\|_{2}+\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}.$
@@
    $\|{\nabla L(\beta^{*})}\|_{2}\leq C_{1}\frac{\sqrt{\log n}}{\sqrt{n}}.$
    For the second term, we analyze $\nabla(\widetilde{L}-L)(\beta^{*})=\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})$, where
@@
    $\|T_{i}(\beta^{*})\|_{2}\leq|y_{i}|\|{\tilde{x}_{i}-x_{i}}\|_{2}+|\sigma(\beta^{*\top}\tilde{x}_{i})|\|{\tilde{x}_{i}-x_{i}}\|_{2}+|\sigma(\beta^{*\top}\tilde{x}_{i})-\sigma(\beta^{*\top}x_{i})|\|{x_{i}}\|_{2}.$
    Using bounds $|y_{i}|\leq 1$, $|\sigma(\cdot)|\leq 1$, $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta$ (which is Assumption C4), and the $\frac{1}{4}$-Lipschitz property of $\sigma$:
    $\|T_{i}(\beta^{*})\|_{2}\leq 1\cdot\delta+1\cdot\delta+\frac{1}{4}|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\|{x_{i}}\|_{2}.$
    Applying Cauchy-Schwarz to $|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\leq\|{\beta^{*}}\|_{2}\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\|{\beta^{*}}\|_{2}\delta$, under $\mathcal{E}_{0}$ and C2, we have
    $\|T_{i}(\beta^{*})\|_{2}$ $\leq 2\delta+\frac{1}{4}\|{\beta^{*}}\|_{2}\delta\|{x_{i}}\|_{2}=\left(2+\frac{1}{4}\|{\beta^{*}}\|_{2}\|{x_{i}}\|_{2}\right)$
@@
    $=\left(2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)\right)\delta.$
@@
    $\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}$ $=\|{\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})}\|_{2}\leq\frac{1}{n}\sum_{i=1}^{n}\|T_{i}(\beta^{*})\|_{2}=\left(2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)\right)\delta.$
    Let $C_{G}^{*}(n)=2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)=O((\log n)(\log m))$. We have:
    $\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}\leq C_{G}^{*}(n)\delta.$
    Combining with the bound for $\|{\nabla L(\beta^{*})}\|$ on event $\mathcal{E}_{1}$:
    $\rho_{n}:=\|{\nabla\widetilde{L}(\beta^{*})}\|\leq\frac{C_{1}\log n}{\sqrt{n}}+C_{G}^{*}(n)\delta\quad(\text{holds on }\mathcal{E}_{0}\cap\mathcal{E}_{1}).$
@@
    $O\left(\frac{\log n}{\sqrt{n}}\right)+O(\delta(\log n)(\log m))\leq\frac{\mu r_{1}}{8}.$
    Since $\delta=o(\frac{1}{(\log n)(\log m)})$ by C5, the inequality holds for sufficiently large $n$.
  • Injectivity: On $\mathcal{E}_{3}$, Jacobian $J_{\phi}(\beta)=\nabla^{2}\widetilde{L}(\beta)\succeq\tilde{\mu}I\succ 0$ in $B_{r_{1}}(\beta^{*})$, implying $\phi$ is injective on $B_{r_{1}}(\beta^{*})$.
All conditions hold on $\mathcal{E}$ for sufficiently large $n$. Therefore, we can apply Lemma F.3 for Localization. On $\mathcal{E}$, Lemma F.3 guarantees the existence of $\hat{\beta}_{soln}\in B_{r_{1}}(\beta^{*})$, such that $\nabla\widetilde{L}(\hat{\beta}_{soln})=0$. Since $\widetilde{L}(\beta)$ is strictly convex on $B_{r_{1}}(\beta^{*})$ (as $\nabla^{2}\widetilde{L}\succeq\tilde{\mu}I>0$ on $\mathcal{E}_{3}$), $\hat{\beta}_{soln}$ is the unique minimizer in $B_{r_{1}}(\beta^{*})$. Global convexity implies that $\hat{\beta}_{soln}$ is the unique global minimum $\hat{\beta}$. Therefore, on $\mathcal{E}$, we have that $\hat{\beta}$ exists, is unique, and satisfies
$\|\hat{\beta}-\beta^{*}\|_{2}\leq r_{1}.$
Next, we derive a bound on $\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}$. We will first show that $\hat{\beta}^{\,(true)}$ must also be in $B_{r_{1}}(\beta^{*})$, under $\mathcal{E}$. Recall that that
$\nabla^{2}L(\beta)\;\succeq\;\frac{\mu}{4}\,I\quad\text{for all }\|\beta-\beta^{*}\|\leq r_{1},\quad\text{and}\quad\|\nabla L(\beta^{*})\|_{2}\;\leq\;C_{1}\sqrt{\frac{\log n}{n}}\,.$
@@
$\nabla^{2}L(\beta)\;\succeq\;\frac{\mu}{4}\,I\quad\forall\;\|\beta-\beta^{*}\|\leq r_{1},\qquad\|\nabla L(\beta^{*})\|_{2}\;\leq\;\frac{\mu}{8}\,r_{1}.$
@@
$L(\beta_{0})=L(\beta^{*})+\nabla L(\beta^{*})^{\top}(\beta_{0}-\beta^{*})+\frac{1}{2}\,(\beta_{0}-\beta^{*})^{\top}\Bigl{[}\int_{0}^{1}\nabla^{2}L\bigl{(}\beta^{*}+s(\beta_{0}-\beta^{*})\bigr{)}\,ds\Bigr{]}\,(\beta_{0}-\beta^{*}).$
On the event $\mathcal{E}_{2}$, for every point $\xi$ on that segment we have $\nabla^{2}L(\xi)\succeq\frac{\mu}{4}\,I$, so the averaged Hessian $H_{\rm avg}=\int_{0}^{1}\nabla^{2}L(\beta^{*}+s(\beta_{0}-\beta^{*}))\,ds$ also satisfies $H_{\rm avg}\succeq\frac{\mu}{4}\,I$. Hence
$L(\beta_{0})-L(\beta^{*})\;\geq\;-\|\nabla L(\beta^{*})\|_{2}\,r_{1}\;+\;\tfrac{1}{2}\cdot\tfrac{\mu}{4}\,r_{1}^{2}\;=\;\frac{\mu}{8}\,r_{1}^{2}\;-\;\|\nabla L(\beta^{*})\|_{2}\,r_{1}.$
Since on $\mathcal{E}_{1}$ we have $\|\nabla L(\beta^{*})\|_{2}\leq\frac{\mu}{8}\,r_{1}$, it follows that
$L(\beta_{0})\;\geq\;L(\beta^{*}).$
@@
Notice $w$ and $(\beta_{0}-\beta^{*})$ point in exactly the same direction.
Since $L$ is convex and differentiable everywhere, we have
$L(\beta)=L(\beta_{0}+w)\;\geq\;L(\beta_{0})\;+\;\nabla L(\beta_{0})^{\top}w.$
@@
$\nabla L(\beta_{0})^{\top}w\;=\;(\|\beta-\beta^{*}\|-r_{1})\;v^{\top}\nabla L(\beta_{0})\;>\;0.$
@@
$L(\beta)\;\geq\;L(\beta_{0})\;+\;\nabla L(\beta_{0})^{\top}w\;\;>\;\;L(\beta_{0}).$
@@
Now, since $\hat{\beta}^{(\mathrm{true})}\in B_{r_{1}}(\beta^{*})$ and $\nabla L(\hat{\beta}^{(\mathrm{true})})=0$, using Taylor expansion of the gradient around $\beta^{*}$, we have:
$0=\nabla L\bigl{(}\hat{\beta}^{(\mathrm{true})}\bigr{)}=\nabla L(\beta^{*})\;+\;\underbrace{\int_{0}^{1}\nabla^{2}L\bigl{(}\beta^{*}+t(\hat{\beta}^{(\mathrm{true})}-\beta^{*})\bigr{)}\,dt}_{=:H_{\rm avg}}\;\bigl{(}\hat{\beta}^{(\mathrm{true})}-\beta^{*}\bigr{)}.$
By construction $H_{\rm avg}\succeq\frac{\mu}{4}I$. Therefore
$\frac{\mu}{4}\;\bigl{\|}\hat{\beta}^{(\mathrm{true})}-\beta^{*}\bigr{\|}_{2}\;\leq\;\bigl{\|}H_{\rm avg}\,(\hat{\beta}^{(\mathrm{true})}-\beta^{*})\bigr{\|}_{2}=\|\nabla L(\beta^{*})\|_{2}\;\leq\;C_{0}\sqrt{\frac{\log n}{n}}.$
Next, we get the bound for $\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}$. Let $\Delta=\hat{\beta}-\hat{\beta}^{\,(true)}$. On $\mathcal{E}$, $\hat{\beta},\hat{\beta}^{\,(true)}\in B_{r_{1}}(\beta^{*})$, so the segment $\hat{\beta}^{\,(true)},\hat{\beta}]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}\subset\mathcal{E}$, $\nabla^{2}\widetilde{L}(\beta)\succeq\tilde{\mu}I$ on this segment. Let $(\Delta^{\prime})=\widetilde{L}(\hat{\beta}^{\,(true)}+\Delta^{\prime}). ff$ is $\tilde{\mu}-$strongly convex along $0,\Delta]$. Gradient monotonicity gives $\langle\nabla f(\Delta)-\nabla f(0),\Delta-0\rangle\geq\tilde{\mu}\|\Delta\|_{2}^{2}$. With $\nabla f(\Delta)=0$ and $\nabla f(0)=\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})$:
$\langle-\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)}),\Delta\rangle\geq\tilde{\mu}\|\Delta\|_{2}^{2}.$
By Cauchy-Schwarz:
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\|\Delta\|_{2}\geq\tilde{\mu}\|\Delta\|_{2}^{2}.$
If $\|\Delta\|_{2}\neq 0$:
$\|\Delta\|_{2}\leq\frac{1}{\tilde{\mu}}\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}.$
We use the similar derivations we had before to bound the gradient differences. Let $B(n)=\|\hat{\beta}^{\,(true)}\|_{2}=O(\log n)$ on $\mathcal{E}$. Then
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq\left(2+O((\log n)(\log m))\right)\delta.$
Let $C_{G}(n)=(2+O((\log n)(\log m)))$. Then $\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq C_{G}(n)\delta$. Substituting:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}\leq\frac{C_{G}(n)}{\tilde{\mu}}\delta=\frac{8C_{G}(n)}{\mu}\delta.$
Let $C_{\Delta}(n)=\frac{8C_{G}(n)}{\mu}=O((\log n)(\log m))$. Then on $\mathcal{E}$:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}\leq C_{\Delta}(n)\delta=O(\delta(\log n)(\log m)).$
@@
$\|\hat{\beta}-\beta^{*}\|_{2}$ $\leq\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}+\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}$
$\leq O\left(\max\left\{\delta(\log n)(\log m),\frac{\log n}{\sqrt{n}}\right\}\right).$

### v3

@@ D.1 Consistency under inner product latent space models
  • $\mathcal{E}_{0}$: Event where $\|x_{i}\|\leq C_{R}\log n=:R^{\prime}(n)$ for all $i=1,\ldots,n$. $\mathbb{P}(\mathcal{E}_{0})\geq 1-n^{-(k+1)}$.
  • $\mathcal{E}_{1}$: Event where $\|\nabla\mathcal{L}(\beta^{*})\|\leq C_{1}\sqrt{(\log n)/n}$. From Lemma D.1, $\mathbb{P}(\mathcal{E}_{1})\geq 1-c_{1}n^{-(k+1)}$.
  • $\mathcal{E}_{2}$: $\nabla^{2}\mathcal{L}(\beta)\succeq(\mu/4)I_{d}$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. From Lemma D.2, $\mathbb{P}(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$.
  • $\mathcal{E}_{3}$: $\nabla^{2}\tilde{\mathcal{L}}(\beta)\succeq(\mu/8)I_{d}=:\tilde{\mu}I_{d}$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$.
@@
$\|\nabla^{2}\tilde{\mathcal{L}}(\beta)-\nabla^{2}\mathcal{L}(\beta)\|_{\rm op}\leq K\delta,$
where $K$ depends only on the $\|x_{i}\|$‐bounds (and thus is $\mathcal{O}((\log n)^{2})$ on $\mathcal{E}_{0}$). Hence
$\nabla^{2}\tilde{\mathcal{L}}(\beta)\succeq\nabla^{2}\mathcal{L}(\beta)-K\delta I_{d}\succeq\frac{\mu}{4}I_{d}-K\delta I_{d}.$
@@
$\nabla^{2}\tilde{\mathcal{L}}(\beta)\succeq\frac{\mu}{8}I_{d},\quad\forall\|\beta-\beta^{*}\|\leq r_{1},$
@@
$\mathcal{E}_{2}\subset\mathcal{E}_{3}\implies\mathbb{P}(\mathcal{E}_{3})\geq\mathbb{P}(\mathcal{E}_{2}).$
Let $\mathcal{E}=\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$, and we have $\mathbb{P}(\mathcal{E})\geq 1-c_{5}n^{-(k+1)}$.
Next, we work under the conditions from $\mathcal{E}$. We want to use Lemma F.3 to show the unique solution. Set $\phi(\beta)=\nabla\tilde{\mathcal{L}}(\beta)$, $x^{*}=\beta^{*}$, $y^{*}=\nabla\tilde{\mathcal{L}}(\beta^{*})$, $y=0$, $\delta_{1}=r_{1}$.
@@
  • Lower bound: For $\|{\beta-\beta^{*}}\|=r_{1}$. Let $H_{\mathrm{avg}}(\beta)=\int_{0}^{1}\nabla^{2}\tilde{\mathcal{L}}(\beta^{*}+t(\beta-\beta^{*}))\,\mathrm{d}t$. The segment $\beta^{*},\beta]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}$, $\nabla^{2}\tilde{\mathcal{L}}(\cdot)\succeq\tilde{\mu}I_{d}$ in $B_{r_{1}}(\beta^{*})$, thus $H_{\mathrm{avg}}(\beta)\succeq\tilde{\mu}I_{d}$. Then, $\|{\phi(\beta)-y^{*}}\|=\|{H_{\mathrm{avg}}(\beta)(\beta-\beta^{*})}\|\geq\lambda_{\min}(H_{\mathrm{avg}}(\beta))\|{\beta-\beta^{*}}\|\geq\tilde{\mu}r_{1}$. Set $\rho=\tilde{\mu}r_{1}$. The condition $\min_{\|{\beta-\beta^{*}}\|=r_{1}}\|{\phi(\beta)-y^{*}}\|\geq\rho$ holds on $\mathcal{E}_{3}$.
  • Upper bound: We need to show $\|{y-y^{*}}\|=\|{0-\nabla\tilde{\mathcal{L}}(\beta^{*})}\|=\|{\nabla\tilde{\mathcal{L}}(\beta^{*})}\|\leq\rho=\tilde{\mu}r_{1}$. We decompose the gradient using the triangle inequality:
    $\|{\nabla\tilde{\mathcal{L}}(\beta^{*})}\|$ $=\|{\nabla\mathcal{L}(\beta^{*})+\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta^{*})}\|$
    $\leq\|{\nabla\mathcal{L}(\beta^{*})}\|+\|{\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta^{*})}\|.$
@@
    $\|{\nabla\mathcal{L}(\beta^{*})}\|\leq C_{1}\sqrt{\frac{\log n}{n}}.$
    For the second term, we analyze $\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta^{*})=\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})$, where
@@
    $\|T_{i}(\beta^{*})\|\leq|y_{i}|\|{\tilde{x}_{i}-x_{i}}\|+|\sigma(\beta^{*\top}\tilde{x}_{i})|\|{\tilde{x}_{i}-x_{i}}\|+|\sigma(\beta^{*\top}\tilde{x}_{i})-\sigma(\beta^{*\top}x_{i})|\|{x_{i}}\|.$
    Using bounds $|y_{i}|\leq 1$, $|\sigma(\cdot)|\leq 1$, $\|{\tilde{x}_{i}-x_{i}}\|\leq\delta$ (which is Assumption C4), and the $\frac{1}{4}$-Lipschitz property of $\sigma$:
    $\|T_{i}(\beta^{*})\|\leq 1\cdot\delta+1\cdot\delta+\frac{1}{4}|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\|{x_{i}}\|.$
    Applying Cauchy-Schwarz inequality to $|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\leq\|{\beta^{*}}\|\|{\tilde{x}_{i}-x_{i}}\|\leq\|{\beta^{*}}\|\delta$, under $\mathcal{E}_{0}$ and Assumption C2, we have
    $\|T_{i}(\beta^{*})\|$ $\leq 2\delta+\frac{1}{4}\|{\beta^{*}}\|\delta\|{x_{i}}\|$
    $=\left(2+\frac{1}{4}\|{\beta^{*}}\|\|{x_{i}}\|\right)\delta$
@@
    $=\left(2+\frac{C_{\beta}C_{R}}{4}\log n\log m\right)\delta.$
@@
    $\|{\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta^{*})}\|$ $=\left\|\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})\right\|\leq\frac{1}{n}\sum_{i=1}^{n}\|T_{i}(\beta^{*})\|=\left(2+\frac{C_{\beta}C_{R}}{4}\log n\log m\right)\delta.$
    Let $C_{G}^{*}(n)=2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)=\mathcal{O}(\log n\log m)$. We have:
    $\|{\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta^{*})}\|_{2}\leq C_{G}^{*}(n)\delta.$
    Combining with the bound for $\|{\nabla\mathcal{L}(\beta^{*})}\|$ on event $\mathcal{E}_{1}$:
    $\rho_{n}:=\|{\nabla\tilde{\mathcal{L}}(\beta^{*})}\|\leq\frac{C_{1}\log n}{\sqrt{n}}+C_{G}^{*}(n)\delta\quad(\text{holds on }\mathcal{E}_{0}\cap\mathcal{E}_{1}).$
@@
    $\mathcal{O}\left(\frac{\log n}{\sqrt{n}}\right)+\mathcal{O}(\delta\log n\log m)\leq\frac{\mu r_{1}}{8}.$
    Since $\delta=o(1/(\log n\log m))$ by Assumption C5, the inequality holds for sufficiently large $n$.
  • Injectivity: On $\mathcal{E}_{3}$, Jacobian $J_{\phi}(\beta)=\nabla^{2}\tilde{\mathcal{L}}(\beta)\succeq\tilde{\mu}I_{d}\succ 0$ in $B_{r_{1}}(\beta^{*})$, implying $\phi$ is injective on $B_{r_{1}}(\beta^{*})$.
All conditions hold on $\mathcal{E}$ for sufficiently large $n$. Therefore, we can apply Lemma F.3 for Localization. On $\mathcal{E}$, Lemma F.3 guarantees the existence of $\hat{\beta}_{\mathrm{soln}}\in B_{r_{1}}(\beta^{*})$, such that $\nabla\tilde{\mathcal{L}}(\hat{\beta}_{\mathrm{soln}})=0$. Since $\tilde{\mathcal{L}}(\beta)$ is strictly convex on $B_{r_{1}}(\beta^{*})$ (as $\nabla^{2}\tilde{\mathcal{L}}\succeq\tilde{\mu}I_{d}>0$ on $\mathcal{E}_{3}$), $\hat{\beta}_{\mathrm{soln}}$ is the unique minimizer in $B_{r_{1}}(\beta^{*})$. Global convexity implies that $\hat{\beta}_{\mathrm{soln}}$ is the unique global minimum $\hat{\beta}$. Therefore, on $\mathcal{E}$, we have that $\hat{\beta}$ exists, is unique, and satisfies
$\|\hat{\beta}-\beta^{*}\|\leq r_{1}.$
Next, we derive a bound on $\|\hat{\beta}_{\rm true}-\beta^{*}\|$. We will first show that $\hat{\beta}_{\rm true}$ must also be in $B_{r_{1}}(\beta^{*})$, under $\mathcal{E}$. Recall that that
$\nabla^{2}\mathcal{L}(\beta)\succeq\frac{\mu}{4}I_{d},\quad\forall\|\beta-\beta^{*}\|\leq r_{1},\quad\text{and}\quad\|\nabla\mathcal{L}(\beta^{*})\|\leq C_{1}\sqrt{\frac{\log n}{n}}.$
@@
$\nabla^{2}\mathcal{L}(\beta)\succeq\frac{\mu}{4}I_{d},\quad\forall\|\beta-\beta^{*}\|\leq r_{1},\quad\text{and}\quad\|\nabla\mathcal{L}(\beta^{*})\|\leq\frac{\mu}{8}r_{1}.$
@@
$\mathcal{L}(\beta_{0})=\mathcal{L}(\beta^{*})+\nabla\mathcal{L}(\beta^{*})^{\top}(\beta_{0}-\beta^{*})+\frac{1}{2}(\beta_{0}-\beta^{*})^{\top}\left[\int_{0}^{1}\nabla^{2}\mathcal{L}(\beta^{*}+s(\beta_{0}-\beta^{*}))\,\mathrm{d}s\right](\beta_{0}-\beta^{*}).$
On the event $\mathcal{E}_{2}$, for every point $\xi$ on that segment we have $\nabla^{2}\mathcal{L}(\xi)\succeq(\mu/4)I_{d}$, so the averaged Hessian $H_{\rm avg}=\int_{0}^{1}\nabla^{2}\mathcal{L}(\beta^{*}+s(\beta_{0}-\beta^{*}))\,\mathrm{d}s$ also satisfies $H_{\rm avg}\succeq(\mu/4)I_{d}$. Hence
$\mathcal{L}(\beta_{0})-\mathcal{L}(\beta^{*})\geq-\|\nabla\mathcal{L}(\beta^{*})\|r_{1}+\frac{1}{2}\cdot\frac{\mu}{4}r_{1}^{2}=\frac{\mu}{8}r_{1}^{2}-\|\nabla\mathcal{L}(\beta^{*})\|r_{1}.$
Since on $\mathcal{E}_{1}$ we have $\|\nabla\mathcal{L}(\beta^{*})\|\leq\mu r_{1}/8$, it follows that
$\mathcal{L}(\beta_{0})\geq\mathcal{L}(\beta^{*}).$
@@
Notice $w$ and $\beta_{0}-\beta^{*}$ point in exactly the same direction.
Since $\mathcal{L}$ is convex and differentiable everywhere, we have
$\mathcal{L}(\beta)=\mathcal{L}(\beta_{0}+w)\geq\mathcal{L}(\beta_{0})+\nabla\mathcal{L}(\beta_{0})^{\top}w.$
@@
$\nabla\mathcal{L}(\beta_{0})^{\top}w=(\|\beta-\beta^{*}\|-r_{1})v^{\top}\nabla\mathcal{L}(\beta_{0})>0.$
@@
$\mathcal{L}(\beta)\geq\mathcal{L}(\beta_{0})+\nabla\mathcal{L}(\beta_{0})^{\top}w>\mathcal{L}(\beta_{0}).$
@@
Now, since $\hat{\beta}_{\rm true}\in B_{r_{1}}(\beta^{*})$ and $\nabla\mathcal{L}(\hat{\beta}_{\rm true})=0$, using Taylor expansion of the gradient around $\beta^{*}$, we have:
$0=\nabla\mathcal{L}\bigl(\hat{\beta}_{\rm true}\bigr)=\nabla\mathcal{L}(\beta^{*})+\underbrace{\int_{0}^{1}\nabla^{2}\mathcal{L}(\beta^{*}+t(\hat{\beta}_{\rm true}-\beta^{*}))\,\mathrm{d}t}_{=:H_{\rm avg}}\;(\hat{\beta}_{\rm true}-\beta^{*}).$
By construction $H_{\rm avg}\succeq(\mu/4)I_{d}$. Therefore
$\frac{\mu}{4}\|\hat{\beta}_{\rm true}-\beta^{*}\|\leq\|H_{\rm avg}(\hat{\beta}_{\rm true}-\beta^{*})\|=\|\nabla\mathcal{L}(\beta^{*})\|\leq C_{0}\sqrt{\frac{\log n}{n}}.$
Next, we get the bound for $\|\hat{\beta}-\hat{\beta}_{\rm true}\|$. Let $\Delta=\hat{\beta}-\hat{\beta}_{\rm true}$. On $\mathcal{E}$, $\hat{\beta},\hat{\beta}_{\rm true}\in B_{r_{1}}(\beta^{*})$, so the segment $\hat{\beta}_{\rm true},\hat{\beta}]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}\subset\mathcal{E}$, $\nabla^{2}\tilde{\mathcal{L}}(\beta)\succeq\tilde{\mu}I_{d}$ on this segment. Let $(\Delta^{\prime})=\tilde{\mathcal{L}}(\hat{\beta}_{\rm true}+\Delta^{\prime}). ff$ is $\tilde{\mu}-$strongly convex along $0,\Delta]$. Gradient monotonicity gives $\langle\nabla f(\Delta)-\nabla f(0),\Delta-0\rangle\geq\tilde{\mu}\|\Delta\|^{2}$. With $\nabla f(\Delta)=0$ and $\nabla f(0)=\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})$:
$\langle-\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true}),\Delta\rangle\geq\tilde{\mu}\|\Delta\|^{2}.$
By Cauchy-Schwarz inequality:
$\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|\|\Delta\|\geq\tilde{\mu}\|\Delta\|^{2}.$
If $\|\Delta\|\neq 0$:
$\|\Delta\|\leq\frac{1}{\tilde{\mu}}\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|.$
We use the similar derivations we had before to bound the gradient differences. Let $B(n)=\|\hat{\beta}_{\rm true}\|=\mathcal{O}(\log n)$ on $\mathcal{E}$. Then
$\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|\leq(2+\mathcal{O}(\log n\log m))\delta.$
Let $C_{G}(n)=2+\mathcal{O}(\log n\log m)$. Then $\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|\leq C_{G}(n)\delta$. Substituting:
$\|\hat{\beta}-\hat{\beta}_{\rm true}\|\leq\frac{C_{G}(n)}{\tilde{\mu}}\delta=\frac{8C_{G}(n)}{\mu}\delta.$
Let $C_{\Delta}(n)=8C_{G}(n)/\mu=\mathcal{O}(\log n\log m)$. Then on $\mathcal{E}$:
$\|\hat{\beta}-\hat{\beta}_{\rm true}\|\leq C_{\Delta}(n)\delta=\mathcal{O}(\delta\log n\log m).$
@@
$\|\hat{\beta}-\beta^{*}\|$ $\leq\|\hat{\beta}-\hat{\beta}_{\rm true}\|+\|\hat{\beta}_{\rm true}-\beta^{*}\|$
$\leq\mathcal{O}\left(\max\left\{\delta\log n\log m,\frac{\log n}{\sqrt{n}}\right\}\right).$